A plastic sphere floats in water with 50.0 % of it
Answer #1

weight of water displaced=weight of sphere, so water=2 sphere, so sphere's density = 1/2 of water's

weight of another fluid displaced = weight of sphere, so fluid= 3.33 sphere,

so sphere's density = 0.3 of fluid's = 1/2 of water's.

sphere's density = 500 kg/m^3

So fluid's density = 10/3 * 1/2 water = 1.667 water's = 1667 kg/m^3